OpenStudy (rational):

how do i explain this to a 10th grader \[\arccos(x)+\arccos(-x)=\pi\]

OpenStudy (rational):

I made him see that \(\cos(\pi-x)=\cos(x)\) on unit circle but he has difficulty comprehending the inverse functions

OpenStudy (dan815):